# Knowledge of Pretrained Language Models on Surface Information of Tokens

### 15 Feb 2024 v1

@@ 4.1 Examination Method
Here, $f_{\mathrm{length}}(\cdot)$ is an MLP [^7] For all MLPs used in this study, the number of layers was three and the activation function was ReLU. Additionally, the embeddings of subwords/words were frozen during the training of MLPs.. We trained the MLP parameters to minimize the mean squared error (MSE) between the predicted value and the true number of characters.
@@ 4.2 Results
In the word-level experiments, larger models were not clearly superior to smaller models. These results imply that the encoders of PLMs are not capable of handling the length information stored in subword-level embeddings. Figure 3 presents a box plot comparing the predicted and actual lengths of inputs from the word-level experiment with BERT-base-cased, which has the lowest scores in Table 2. One can see that the encoded word embeddings retained accurate length relationships, even though the exact word lengths could not be reconstructed from the embeddings. Additionally, one can see that a longer surface makes it more difficult to predict a correct length. This may be because most of the input tokens used to train the MLPs were relatively short (approximately two to seven characters).
@@
Figure 3: Comparison of predicted lengths and true lengths of word-level inputs in the BERT-base-case. The red line indicates correct predictions.
@@ 8 Conclusion
This study was limited to two languages (English and Japanese) with different writing system characteristics. Experiments with languages with right-to-left writing systems such as Arabic and other languages with low-resource corpora could lead to different conclusions. Additionally, our experiments focused on LLMs with relatively small numbers of parameters (up to 7B) in consideration of limited computational resources. Although this shortcoming could be resolved by using larger numbers of parameters, current state-of-the-art models such ChatGPT seem to be incapable of handing surface information, as shown in Figure 1. We excluded ChatGPT from our experiments because we could not guarantee the reproducibility of output results and because concrete architectures and parameters are not publicly available.

### 22 Feb 2024 v2

@@ 4.1 Examination Method
Here, $f_{\mathrm{length}}(\cdot)$ is an MLP [^7] For all MLPs used in this study, the number of layers was three and the activation function was ReLU. Additionally, the embeddings of subwords/words were frozen during the training of MLPs. The size of hidden parameters was 2,096.. We trained the MLP parameters to minimize the mean squared error (MSE) between the predicted value and the true number of characters.
@@ 4.2 Results
In the word-level experiments, larger models were not clearly superior to smaller models. These results imply that the encoders of PLMs are not capable of handling the length information stored in subword-level embeddings. Figure 3 presents a box plot comparing the predicted and actual lengths of inputs from the word-level experiment with BERT-base-cased, which has the lowest scores in Table 2. One can see that the encoded word embeddings retained almost accurate length relationships, even though the exact word lengths could not be reconstructed from the embeddings. Additionally, one can see that a longer surface makes it more difficult to predict a correct length. This may be because most of the input tokens used to train the MLPs were relatively short (approximately three to seven characters).
@@
Figure 3: Comparison of predicted lengths and true lengths of word-level inputs in the BERT-base-cased. The red line indicates correct predictions.
@@ 8 Conclusion
This study was limited to two languages (English and Japanese) with different writing system characteristics. Experiments with languages with right-to-left writing systems such as Arabic and other languages with low-resource corpora could lead to different conclusions. Additionally, our experiments focused on LLMs with relatively small numbers of parameters (up to 7B) in consideration of limited computational resources. Although this shortcoming could be resolved by using larger numbers of parameters, current state-of-the-art models such ChatGPT seem to be incapable of handing surface information, as shown in Figure 1. We excluded models which are accessible only via API such as ChatGPT from our experiments because we could not guarantee the reproducibility of output results and because concrete architectures and parameters are not publicly available. All experiments were conducted on NIVIDIA RTX A6000 and each training of the examination used in this paper was less than six hours.